OpenStudy (anonymous):

Sheldon and Leonard are preparing for their annual trip to Comic-Con, a comic book convention. After careful analysis of route options, and based on traffic patterns, Leonard determines the following: while driving at a constant pace, they will travel 50 miles in the first hour, 255 miles in the next 3 hours, and 120 miles in the next 2 hours. Create a piecewise function representing Leonard and Sheldon’s traveling time and distance.

OpenStudy (anonymous):

let t = time speed in first hour = 50 miles / 1 hour = 50 miles per hour = 50 mph speed in next 3 hours = 255 miles / 3 hours = 85 mph speed in next 2 hours = 120/2 = 60 mph distance traveled = 50t when 0 < t ≤ 1 = 50 + 85t when 1 < t ≤ 4 = 50 + 255 + 60t when 4 < t ≤ 6

OpenStudy (anonymous):

so then what is the domain and range?

OpenStudy (anonymous):

you should be able to figure that out with my reply :)